Assume 20% of the population are right handed and the remainder are left handed. If you stop the next 7 persons you meet, what is the probability that all will be left handed, right Handed, and exactly 2 will be left handed. Can you show me examples how to figure this out?
Answer by sudhanshu_kmr(1152)

Probability that a person is right handed = 20/100 = 2/ 10 = 1/5

then, the probability of left handed = 1 - 1/5 = 4/5


CASE 1: All 7 are left handed

Probability = (4/5) * (4/5) * .... 7 times


= (4/5)^7



CASE 2: All 7 are right handed
Probability = (1/5)* (1/5) * (1/5)*....... 7 times
= (1/5)^7


CASE 3: Exactly 2 will be left handed

so, 2 are left handed and 5 are right handed

by using binomial probability


required probability = 7C2 * (4/5)^2 * (1/5)^5



Sometimes there may be typing mistake in solution of a problem, so please ignore it. Understand the concept and try the problem yourself. If there is problem related to concept, contact at sudhanshu.cochin@gmail.com.
Best of luck.......